calculate the mass of urea required in making 2.5kg of 0.25 molal aq. solution

Grade:11

1 Answers

Aashna
6 years ago
m = Moles of solute/mass of solvent in kgMole of solute =. Mass of urea /molecular mass of ureaMoles = m/60m = 0.25mass of solvent 2.5 kg Moles = 2.5*.25= .625.625 = m/60 = .625*60 = 3.75 g